OpenStudy (anonymous):

okay so whos pretty good at scale factoring and demensions? a really need help unedrstanding this assignment please! <3

OpenStudy (anonymous):

@jim_thompson5910 maybe you? please this is really important for me to understand

OpenStudy (anonymous):

post your question and i will try my best?

OpenStudy (anonymous):

alright well its not really a question its more advice and etc. okay so here i go, ill post the directions and ill tell you my thoughts

OpenStudy (anonymous):

ok go on!

OpenStudy (anonymous):

1. Make a list to identify each object in the room that will be scaled, such as doors, windows, walls, and furniture. Use an appropriately shaped polygon to represent each object. For example, use a rectangle to represent a bed or couch. A minimum of six objects is required. 2. Determine the scale factor you will use and include the units. 3. Complete a chart, like the one below, to organize your data. Show all calculations and include units. Be sure to specify your scale factor.

OpenStudy (anonymous):

the first one im gonna measure is a door, 6ft by 2ft. but first i need a scale factor which im assuming can be anything right? so lets say every ..... yeah im confused.. help lol

jimthompson5910 (jim_thompson5910):

a nice measurement to work with is an inch

jimthompson5910 (jim_thompson5910):

so if you say that 1 inch = 1 foot, then that is your scale factor

jimthompson5910 (jim_thompson5910):

so your 6 ft by 2 ft door in real life would scale down to 6 in by 2 in on paper

OpenStudy (anonymous):

okay that works, so another question is, on the chart the first part is to identify your object(door), so the demension would be 6ft by 2ft right?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

or you can say instead of 1 in = 1 ft, you can do 1 in = 12 in but the first scale factor is easier to use

jimthompson5910 (jim_thompson5910):

yes assuming you measured it correctly

OpenStudy (anonymous):

okay so if 1in = 1 ft and we have 6 ft by 2 ft itll be 6 in by 2in like you said now that 6 in by 2in would be my dimension of scaled object correct?

OpenStudy (anonymous):

and how would you show a calcuation for that? even possible? lol

jimthompson5910 (jim_thompson5910):

well you can do this 1 in: 1 ft 2*1 in: 2*1 ft 2 in: 2 ft (width of the door) 1 in: 1 ft 6*1 in: 6*1 ft 6 in: 6 ft (height of the door) but yeah there's not much you can do since you can pretty much see the answer

OpenStudy (anonymous):

yes lol, okay another question, my next object is gonna be a picture, lets say it 8in by 10 in. how would i do this one with a scale factor of 1 in = 1 ft

OpenStudy (anonymous):

Wow ur really good at explaining things... @jim_thompson5910

OpenStudy (anonymous):

this helped me a bit...THXS

jimthompson5910 (jim_thompson5910):

now you're going to run into trouble the picture's dimensions are not in feet, so the scale factor is going to have to bend it a bit 8 inches = 8/12 ft 8 inches = 2/3 ft 1 in: 1 ft 2/3*1 in : 2/3*1 ft 2/3 in : 2/3 ft So in real life, the picture has a width of 8 inches or 2/3 ft. The scale drawing will have a width of 2/3 inches

jimthompson5910 (jim_thompson5910):

10 in = 10/12 ft 10 in = 5/6 ft 1 in: 1 ft 5/6*1 in : 5/6*1 ft 5/6 in : 5/6 ft In real life, the picture has a height of 10 inches or 5/6 ft. The scale drawing will have a width of 5/6 inches

jimthompson5910 (jim_thompson5910):

glad I could help you out a bit LizzyLove<3

jimthompson5910 (jim_thompson5910):

and it might be better to convert to decimal form, but it's up to you

OpenStudy (anonymous):

yeah okay so ill do this one myself and you correct me if i make a mistake...

OpenStudy (anonymous):

i have a window 2ft by 4ft, using the scale factor 1in = 1ft, 2in by 4 ft bc 2 in:2 ft and 4in : 4ft so the scaled would be 4in by 2in right?

OpenStudy (anonymous):

im just gonna keep the rest of my measurements simple with no decimals and all more than a foot lol it might make my life alot easier